OpenStudy (anonymous):

The price (in dollars) p and the quantity demanded q are related by the equation: p^2+2q^2=1100.If R is revenue, dR/dt can be expressed by the following equation: dR/dt=A*dp/dt, where A is a function of just q. A = ? Find dR/dt when q=15 and dp/dt=5. dR/dt = The price (in dollars) p and the quantity demanded q are related by the equation: p^2+2q^2=1100.If R is revenue, dR/dt can be expressed by the following equation: dR/dt=A*dp/dt, where A is a function of just q. A = ? Find dR/dt when q=15 and dp/dt=5. dR/dt = @Mathematics

OpenStudy (anonymous):

i got the first part. dont know the second. did you complete all the other webwork questions?

OpenStudy (anonymous):

did you get #13 second part or 14?

OpenStudy (anonymous):

i got the second part, i don't get the first part?

OpenStudy (anonymous):

#14 0.026 ft/min

OpenStudy (anonymous):

i got 14. only 13 2nd part left. do you know? i have all the others if you need help

OpenStudy (anonymous):

13 first part please, second part answer is -26.67

OpenStudy (anonymous):

thats wrong for me. first part is (2q-550/q)

OpenStudy (anonymous):

how did you get your 13 b?

OpenStudy (anonymous):

my q = 20 and dp/dt= 3

OpenStudy (anonymous):

R = pq R' = pq' + p'q q=15 p'=4 you can calculate for p using the first equation that is given p^2+2q^2=1100 you'll get p= sqrt(650) then you find q' by differentiating p^2+2q^2=1100 and sub the numbers in you'll get q'=-3.3 something then plug them all into R' = pq' + p'q and you'll get the answer

OpenStudy (anonymous):

ohhh, that's why

OpenStudy (anonymous):

ok let me try. good explanation

OpenStudy (anonymous):

ok so i got -6 instead of -3.3

OpenStudy (anonymous):

i am lost after that. can you help

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

your number is different, so your answers wont match mine

OpenStudy (anonymous):

after that just plug everything in R' = pq' + p'q and you'll get R'